OpenStudy (anonymous):

xy=3 dy/dt=2, dx/dt=? x=4 I got y(dx/dt)=-8/y. This is wrong. What did I do?

OpenStudy (anonymous):

@Easyaspi314

OpenStudy (anonymous):

y=3/4 it's given in the beginning that xy=3 and you know x is 4 so 4y=3 y=3/4 the d's in dy/dt cancel out so it's y/t=2 or (3/4)/t=2, 2t=3/4, t=3/8

OpenStudy (anonymous):

Okay. Because we usually do a difference process to get these answers.

OpenStudy (anonymous):

So, how does this help me get dx/dt?

OpenStudy (anonymous):

i don't think you need d because whatever its value is, it will be 1 in the end so it's irrelevant \[\frac{ dx }{ dt }\] The value of the expression is x/t which is 4/(3/8) = 32/3

OpenStudy (anonymous):

ohhh. I think I get what you mean!